OpenStudy (angelwings996):

A cylinder contains ice cream. Its radius is 5 inches and it stands 4 inches tall. The Cool Man ice cream shop sells ice cream cones. If a cone is 5 inches tall and has a diameter of 3 inches, how many ice cream scoops can be scooped out of a cylinder?

OpenStudy (angelwings996):

@Compassionate

OpenStudy (compassionate):

[One moment...]

OpenStudy (angelwings996):

That's fine

OpenStudy (compassionate):

Find the volume of the cylinder. Then find the volume of a scoop of ice cream used for a single cone. Divide that by the volume of the cylinder. The question is ambiguous as it does not fully describe the shape of the ice cream scoop used for the cone. Is it a sphere or half a sphere? Is the entire cone filled with ice cream?

OpenStudy (compassionate):

Volume of a cylinder V=(pi)(r^2)(h) =(3.1416)(5)^2(4) =314.16i^3 Volume of a cone V=(1/3)(pi)(r^2)(h) =(1/3)(3.1416)(3/2)^2(5) =11.781i^3 314.16i^3/11.781^3 =25 cones